# Learning to Animate Images from A Few Videos to Portray Delicate Human Actions

### v1

@@ 4.2 Ablation Studies
In Figure A1, we show two additional examples of human action videos generated by Dream Machine, KLING AI, and FLASH. It can be observed that Dream Machine and KLING AI fail to animate these two actions accurately. The generated videos are available on the webpage.
Figure A1: Comparison of human action videos generated by Dream Machine, KLING AI, and FLASH (our method). For the shoot dance action, both Dream Machine and KLING AI produce unrealistic movements that defy physical laws. In the Ice Bucket Challenge action, neither Dream Machine nor KLING AI accurately captures the motion of pouring ice water from the bucket onto the body. In contrast, FLASH successfully generates both actions with a higher fidelity to the real movements, as shown in the last row. Human faces are anonymized for privacy protection.
A2 More Details of FLASH
A2.1 Preliminaries
@@
Temporal Attention Layers. To capture temporal dynamics in videos, temporal attention layers are introduced into the U-Net [27, 9, 21, 20]. In a temporal attention layer, the input features $\bm{F}_{in}\in\mathbb{R}^{N\times h^{\prime}\times w^{\prime}\times c^{\prime}}$ are first reshaped to $\tilde{\bm{F}}_{in}\in\mathbb{R}^{B\times N\times c^{\prime}}$, where $B=h^{\prime}\times w^{\prime}$. Here, the features at different spatial locations are treated as independent samples. Temporal position encoding are then added, and a self-attention layer is applied to transform $\tilde{\bm{F}}_{in}$ into $\tilde{\bm{F}}_{out}\in\mathbb{R}^{B\times N\times c^{\prime}}$. Finally, $\tilde{\bm{F}}_{out}$ is reshaped back to $\bm{F}_{out}\in\mathbb{R}^{N\times h^{\prime}\times w^{\prime}\times c^{\prime}}$ as output features. The temporal attention layer integrates information from different frames for each spatial location, enabling the learning of temporal changes.
@@
Cross-frame Attention Layers. To enhance temporal consistency across generated frames, cross-frame attention layers are utilized to replace self-attention layers [35, 70]. While self-attention layers use features from all frames as key and value, cross-frame attention layers restrict key and value to the features from the first frame or the previous frame. This approach preserves the appearance of objects and background from the first frame or the previous frame, improving temporal consistency in the generated videos.
@@
Noise-Free Frame Conditioning. To preserve the appearance of the reference image in the image animation task, the noise-free latent reference image is integrated into the U-Net input [70, 52]. During training, the first latent frame remains noise-free, while noise is added only to subsequent latent frames throughout the noising process. Specifically, at the noising step $t$, the latent video $\bm{Z}_{t}=\langle\bm{z}_{t}^{i}\rangle_{i=1}^{N}$ is modified to $\check{\bm{Z}}_{t}=\langle\bm{z}_{0}^{1},\bm{z}_{t}^{2},\cdots,\bm{z}_{t}^{N}\rangle$, where $\bm{z}_{t}^{1}$ is replaced by $\bm{z}_{0}^{1}$, which is noise-free. During inference, a sample $\bm{Z}_{T}$ is drawn from $\mathcal{N}(\bm{0},\,I)$, and $\bm{z}_{T}^{1}$ is substituted with $\bm{z}_{0}^{1}=\mathcal{E}(I)$, where $I$ is the user-provided reference image. The modified latent video $\check{\bm{Z}}_{T}=\langle\bm{z}_{0}^{1},\bm{z}_{T}^{2},\cdots,\bm{z}_{T}^{N}\rangle$ is then used for denoising. This technique effectively carries over the features from the first frame to subsequent frames, ensuring that the appearance of the reference image is preserved in the generated video.
@@
FLASH adopts these components as its base video diffusion model and designs a motion alignment module and a detail enhancement decoder on top of it.
@@
A2.2 Motion Alignment Module
Strongly Augmented Videos. To create a strongly augmented version of an original video, we sequentially apply Gaussian blur and random color adjustments to the original video. This process is designed to preserve the original motion while altering the appearance uniformly across all frames.
  • Gaussian blur: A kernel size is randomly selected from a predefined range, as specified in Sec. A3.2. This kernel size is used to apply Gaussian blur consistently to every frame of the original video, ensuring a uniform level of blur throughout.
  • Random color adjustments: After applying Gaussian blur, we randomly adjust the brightness, contrast, saturation, and hue of the video. For each property, an adjustment factor is randomly chosen from its respective predefined range, detailed in Sec. A3.2. The adjustment with the chosen factor is applied uniformly across all frames to maintain consistent color alterations without introducing cross-frame inconsistencies. We implement the random color adjustments using the ColorJitter function in PyTorch.
@@
In LDM, pixel-level details can be distorted when videos are decoded from the latent space, as even slight perturbations within the latent space can lead to noticeable visual artifacts, compromising the intricate details and transition smoothness. To mitigate this issue, we devise the Detail Enhancement Decoder that extends the image decoder $\mathcal{D}$ with additional layers to propagate multi-scale details from the reference image to the generated frames. Given that motion between the first frame and subsequent frames can vary, the Detail Enhancement Decoder is designed with two branches to capture details from both small and large displacements, addressing these two different types of motion.
We define the levels of both the encoder and decoder as $l\in\{0,1,\cdots,L\}$, with $l=0$ representing the pixel space and $l=L$ representing the latent space. At level $l$, we extract the decoder features of the $i$-th decoding frame, denoted as $\bm{h}_{l}^{i}$, and the encoder features of the reference image, denoted as $\bm{g}_{l}^{1}$. We interpolate $\bm{g}_{l}^{1}$ to match the spatial size of $\bm{h}_{l}^{i}$ and use a fully connected layer to adjust $\bm{g}_{l}^{1}$ to the same number of channels as $\bm{h}_{l}^{i}$, resulting $\tilde{\bm{g}}_{k}^{1}$ as the input of the following two branches, as illustrated in Figure 2 (c) of the main paper.
Warping Branch. This branch aims to retrieve details from nearby areas in $\tilde{\bm{g}}_{l}^{1}$ for each spatial position in $\bm{h}_{l}^{i}$. It takes the channel-wise concatenation of $\bm{h}_{l}^{i}$ and $\tilde{\bm{g}}_{l}^{1}$ as input and applies four convolution layers to estimate motion displacements from $\bm{h}_{l}^{i}$ to $\tilde{\bm{g}}_{l}^{1}$. These displacements determine the sampling positions in $\tilde{\bm{g}}_{l}^{1}$. By warping $\tilde{\bm{g}}_{l}^{1}$ based on the sampling positions, the branch outputs $\hat{\bm{g}}_{l}^{1}$.
Patch Attention Branch. This branch retrieves details from the global scope of $\tilde{\bm{g}}_{l}^{1}$, complementing the local retrieval in the warping branch. It begins by dividing both $\bm{h}_{l}^{i}$ and $\tilde{\bm{g}}_{l}^{1}$ into patches and transforming each patch into features through a fully connected layer. A cross-attention layer is then applied, using the patch features of $\bm{h}_{l}^{i}$ as the query and the patch features of $\tilde{\bm{g}}_{l}^{1}$ as the key and value. The output features $\check{\bm{g}}_{l}^{1}$ is a weighted patch combination of $\tilde{\bm{g}}_{l}^{1}$.
Feature Fusion. To control the amount of detail added to $\bm{h}_{l}^{i}$, a two-layer convolution network is used to learn the fusion weights. The network takes the channel-wise concatenation of $\bm{h}_{l}^{i}$ and $\tilde{\bm{g}}_{l}^{1}$ as input and outputs the fusion weights $\bm{w}^{i}_{l}$, which has the same spatial size as $\bm{h}_{l}^{i}$. The fusion is then performed as:
$\tilde{\bm{h}}_{l}^{i}=\bm{h}_{l}^{i}+\bm{w}^{i}_{l}\odot(\hat{\bm{g}}_{l}^{1}+\check{\bm{g}}_{l}^{1}).$ (6)
The resulting feature $\tilde{\bm{h}}_{l}^{i}$ is then passed to level $l-1$. The details in the generated frames are enhanced through the hierarchical detail propagation in each level.
Learning to Reconstruct Distorted Videos. We train the Detail Enhancement Decoder to retrieve proper details through reconstructing distorted videos to match their ground-truth versions. During training, we first extract $\bm{g}_{l}^{1}$ using the first frame of a training video. Next, we intentionally distort the video using random Gaussian blur, random color adjustments on 80% of the randomly selected regions, and random elastic transformations. The distorted video is then encoded into a latent video. The decoder is trained to reconstruct the ground-truth video from the distorted latent video and the extracted feature $\bm{g}_{l}^{1}$ of the first frame, using MSE loss. This approach encourages the decoder to retrieve relevant details from the reference image to compensate for distorted details in the latent videos.
@@
The details of video distortions are as follows: The random Gaussian blur and random color adjustments follow the implementations described in Sec. A2.2. However, the random color adjustments here differ in that they are applied to only 80% of randomly selected regions rather than to all regions. This modification is intentional, as the goal is to create distorted videos that simulate the perturbations in latent videos, rather than to maintain consistent color changes as in Sec. A2.2. For random elastic transformations, displacement vectors are generated for all pixels based on random offsets sampled from a predefined range (detailed in Sec. A3.2) and are then used to transform each pixel accordingly. We implement the random elastic transformations using the ElasticTransform function in PyTorch.
@@
We conduct experiments on 12 actions selected from the HAA500 dataset [7], which contains 500 human-centric atomic actions, each consisting of 20 short videos. The selected actions include single-person actions (push-up, arm wave, shoot dance, running in place, and sprint run), human-object interactions (soccer shoot, drinking from a cup, balance beam jump, canoeing sprint, chopping wood, and ice bucket challenge), and human-human interactions (hugging human).
@@
Similarity between videos in the same action class. Videos within the same action class do not necessarily share similar visual characteristics, such as scenes, viewing angles, or shot types (e.g., close-up or wide shot), as shown in the examples in Figure A3.
Testing images. For each selected action, we use the first frames from the 4 testing videos as testing images. Additionally, we search online for 2 human images depicting a person beginning the desired action as additional testing images.
Figure A3: Similarity between videos in the same action class. The first row presents three videos depicting the action canoeing sprint, and the second row showcases three videos illustrating the action push-up.
Figure A4: AMT task interface.
Figure A5: User preference rates (%) of different methods.
@@
We use AnimteDiff [22] as the base model, initializing all parameters with its pretrained weights. The spatial resolution is set to $512\times 512$, and the video length is set to 16 frames.
Training of U-Net. We combine features from the first and current frames as keys and values in the cross-frame attention layers. As in [70, 52], we apply noise-free first-frame conditioning (see Supplementary Material Sec. A2.1 for details). Following [30, 46], we redefine the sampling probability distribution to prioritize earlier denoising stages. In the motion alignment modules, we set $\tau$ to 90 and apply motion feature alignment after each temporal attention layer in the U-Net. Inter-frame correspondence relation alignment is applied to 50% of the cross-frame attention layers, selected randomly. For simplicity, we replace $\bm{Q}$ and $\bm{K}$ of the augmented video with those of the original video when calculating $S$, instead of directly replacing $S$. Gaussian blur is applied with a randomly sampled kernel size between 3 and 10. Color adjustments modify brightness, saturation, and contrast by random factors between 0.5 and 1.5, and modify hue by a random factor between -0.25 and 0.25. Before applying strong augmentations to the original video, we first perform random horizontal flipping and random cropping on the original video. We only train the temporal attention layers and the key and value projection matrices of other attention layers. The learning rate is set to $5.0\times 10^{-5}$, with training conducted for 20,000 steps.
Training of Detail Enhancement Decoder. The patch size in the Patch Attention Branch is set to 2. For video distortion, Gaussian blur is applied with a random kernel size between 3 and 10. Color adjustments use random factors for brightness, saturation, and contrast between 0.7 and 1.3, and a random factor for hue between -0.2 and 0.2. Elastic transformation displacement strength is randomly sampled from 1 to 20. We only train the newly added branches, with a learning rate of $1.0\times 10^{-4}$ over 10,000 steps.
@@
In line with previous works [69, 70, 23], we use three CLIP-based metrics to assess text alignment, image alignment, and temporal consistency. (1) Text Alignment: We compute the similarity between each frame and the provided text prompt, averaging the scores across all frames. (2) Image Alignment: Similar to Text Alignment, we replace the text prompt with the provided reference image to compute the image alignment score. (3) Temporal Consistency: We calculate the average similarity between consecutive frame pairs to obtain the temporal consistency score. We use ViT-L/14 from OpenAI [51] for feature extraction. In these three metrics, higher scores indicate better performance.
Following [72], we utilize Fréchet distance to compare generated and real videos. We use CD-FVD [12] to mitigate content bias in the widely used FVD [62]. We use VideoMAE [61], pretrained on SomethingSomethingV2 [17], for feature extraction and calculate distance between real and generated videos. In this metric, lower distances reflect better performance.
To evaluate the similarity between generated and ground-truth videos in the HAA dataset, we calculate the cosine similarity for each pair of the generated and ground-truth videos. (1) Cosine RGB: We extract video features using I3D [3], pretrained on RGB videos, for both the generated and ground truth videos, calculating cosine similarity for each pair. (2) Cosine Flow: We extract optical flow using RAFT [60] and then use I3D [3], pretrained on optical flow data, to extract features for cosine similarity calculation. In these two metrics, higher similarities indicate better performance.
@@
We compare FLASH with several image animation baselines: (1) TI2V-Zero [48], a training-free image animation model based on a pretrained text-to-video model. (2) SparseCtrl [20], a model trained on large-scale datasets that encodes the reference image with a sparse condition encoder and integrates the features into a video generative model. (3) PIA [80], a model trained on large-scale datasets that incorporates the reference image into noisy latent videos. (4) DynamiCrafter [72], a model trained on large-scale datasets that injects the reference image features into generated videos via cross-attention layers and feature concatenation. (5) DreamVideo [68], which adapts subjects and motion using a limited set of samples; we customize motion for each action using the same training videos as FLASH. (6) LAMP [70], which learns motion patterns from a few videos; we train it with the same training videos as our method.
@@
Variant # Videos Per Class joint Training Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)   CD-FVD ($\downarrow$)   Cosine RGB ($\uparrow$)   Cosine Flow ($\uparrow$)
#1 16 22.53 77.10 95.43 1023.30 0.8380 0.6806
#2 16 22.48 76.72 94.91 932.92 0.8398 0.7061
#5 16 22.52 76.35 95.01 906.31 0.8446 0.7224
#1 8 22.70 76.05 94.79 995.43 0.8250 0.6813
#2 8 22.62 74.37 94.40 962.82 0.8330 0.7009
#5 8 22.66 75.02 94.51 943.54 0.8340 0.7201
#1 4 22.22 72.81 94.24 1050.03 0.8140 0.6802
#2 4 22.60 72.00 93.83 1045.49 0.8188 0.7015
#5 4 22.46 72.56 94.22 1031.87 0.8222 0.7183
#5 16 22.52 76.35 95.01 906.31 0.8446 0.7224
#5 16 22.61 77.47 95.39 897.05 0.8501 0.7232
Table A2: Ablation studies on different values of $\tau$ for motion feature alignment, different values of $p$ for inter-frame correspondence relation alignment, and the impact of the warping branch and patch attention branch in the Detail Enhancement Decoder.
Variant     $\tau$     $p$    Warping Branch    Patch Attention Branch Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)   CD-FVD ($\downarrow$)   Cosine RGB ($\uparrow$)   Cosine Flow ($\uparrow$)
#3  90 - - - 22.64 76.48 95.06 920.39 0.8444 0.7140
#3 75 - - - 22.58 76.63 95.16 904.25 0.8438 0.7119
#3 50 - - - 22.57 77.29 95.14 934.84 0.8430 0.7031
#3 25 - - - 22.33 76.52 94.85 930.53 0.8471 0.6979
#4 -  1.0 - - 22.50 76.43 94.91 914.12 0.8422 0.6934
#4 - 0.5 - - 22.70 76.31 94.84 938.21 0.8432 0.7172
#5 90 0.5 22.52 76.35 95.01 906.31 0.8446 0.7224
#6 90 0.5 22.54 76.21 95.35 918.61 0.8463 0.7196
#6 90 0.5 22.71 74.97 95.13 888.05 0.8332 0.7226
#6 90 0.5 22.77 76.22 95.31 908.39 0.8451 0.7233
Figure A6: Failure cases.
@@
We conducted a user study on Amazon Mechanical Turk (AMT), where participants were asked to select the best generated video from a set of candidates. For each action, we randomly selected four different reference images and their corresponding generated videos for this user study. The AMT assessment interface, shown in Figure A4, presented workers with the following instructions: “You will see a reference image on the left and seven human action videos on the right, all generated from that reference image and an action description. Please carefully select the one video in each question that: (1) Best matches the action description and displays the action correctly and smoothly. (2) Maintains the overall appearance of the reference image on the left.” The interface also displayed the reference image and action description.
To identify random clicking, each question was paired with a control question that included a real video of a randomly selected action, alongside clearly incorrect videos, such as static videos or videos from the same action class that did not match the reference image. Main and control questions were randomly shuffled within each question pair, and each pair was evaluated by 10 different workers. Responses from workers who failed the control questions were marked as invalid.
In total, we collected 366 valid responses. The preference rates for different methods are shown in the pie chart in Figure A5. FLASH was preferred in 67% of valid responses, significantly outperforming the next best choices, DynamiCrafter(14%) and LAMP (12%).
@@
In line with the main paper, Variant #1 serves as the baseline, excluding both the Motion Alignment Module and the Detail Enhancement Decoder. Variant #2 uses only strongly augmented videos without any alignment technique. Variants #3, #4, and #5 progressively incorporate motion feature alignment, inter-frame correspondence relation alignment, and both, respectively. Lastly, Variant #6 builds upon Variant #5 by incorporating the Detail Enhancement Decoder.
@@
Applicability with Fewer Training Videos. To assess the few-shot learning capability of the Motion Alignment Module, we conduct experiments using 8 and 4 videos randomly sampled from each action class. The results are shown in Table A1. We observe that Variant #5 consistently outperforms Variants #1 and #2 across different numbers of training videos per action class. The results validate that the Motion Alignment Module enhances the quality of animated videos in different few-shot configurations.

### v2

@@ 5 Conclusion
In Figure A1, we present four examples of animated human action videos from Dream Machine [^1] https://lumalabs.ai/dream-machine, KLING AI [^2] https://www.klingai.com/, Wanx AI [^3] https://tongyi.aliyun.com/wanxiang/, and FLASH. The videos are available on the Webpage. Dream Machine, KLING AI and Wanx AI struggle to animate these actions accurately. In the balance beam jump action, Dream Machine and Wanx AI produce unrealistic, physics-defying movements, while KLING AI generates a jump but fails to depict standard jumps on the balance beam. For the soccer shooting action, all three models fail to generate a correct shooting motion, with the person never kicking the ball. In the shoot dance action, Dream Machine and KLING AI generate unnatural, physically implausible movements, whereas Wanx AI produces dance movements but does not capture the shoot dance correctly. In the Ice Bucket Challenge action, none of the three models accurately portray the motion of pouring ice water from the bucket onto the body. In contrast, FLASH generates these actions with higher fidelity to the real actions.
Figure A1: Comparison of human action videos generated by Dream Machine, KLING AI, Wanx AI and FLASH (our method). Human faces are anonymized for privacy protection.
A2 FLASH
A2.1 Components in Latent Video Diffusion Models
@@
Temporal Attention Layers. To capture temporal dynamics in videos, [33, 11, 26, 25] add temporal attention layers after each spatial attention layer of U-Net. In each temporal attention layer, we first reshape the input features $\bm{F}_{in}\in\mathbb{R}^{N\times h^{\prime}\times w^{\prime}\times c^{\prime}}$ to $\tilde{\bm{F}}_{in}\in\mathbb{R}^{B\times N\times c^{\prime}}$, where $B=h^{\prime}\times w^{\prime}$. Here, we treat the features at different spatial locations as independent samples. Then, we add temporal position encoding to $\tilde{\bm{F}}_{in}$ and employ a self-attention layer to transform $\tilde{\bm{F}}_{in}$ into $\tilde{\bm{F}}_{out}\in\mathbb{R}^{B\times N\times c^{\prime}}$. Finally, we reshape $\tilde{\bm{F}}_{out}$ to $\bm{F}_{out}\in\mathbb{R}^{N\times h^{\prime}\times w^{\prime}\times c^{\prime}}$ as the output features. The temporal attention layer integrates information from different frames for each spatial location, enabling the learning of temporal changes.
@@
Cross-frame Attention Layers. To enhance temporal consistency across generated frames, [43, 92] replace spatial self-attention layers with spatial cross-frame attention layers. While self-attention layers use features from the current frame as key and value, cross-frame attention layers restrict key and value to the features from the first frame. These layers carry over the appearance features from the first frame to subsequent frames, improving temporal consistency in the generated videos.
@@
Noise-Free Frame Conditioning. To further preserve the appearance of the reference image in the image animation task, [92, 68] keep the latent reference image noise-free in the noised latent video. Specifically, at the noising step $t$, the latent video $\bm{Z}_{t}=\langle\bm{z}_{t}^{i}\rangle_{i=1}^{N}$ is modified to $\check{\bm{Z}}_{t}=\langle\bm{z}_{0}^{1},\bm{z}_{t}^{2},\cdots,\bm{z}_{t}^{N}\rangle$, where $\bm{z}_{t}^{1}$ is replaced by $\bm{z}_{0}^{1}$, which is noise-free. During inference, a sample $\bm{Z}_{T}$ is drawn from $\mathcal{N}(\bm{0},\,I)$, and $\bm{z}_{T}^{1}$ is substituted with $\bm{z}_{0}^{1}=\mathcal{E}(I)$, where $I$ is the user-provided reference image. The modified latent video $\check{\bm{Z}}_{T}=\langle\bm{z}_{0}^{1},\bm{z}_{T}^{2},\cdots,\bm{z}_{T}^{N}\rangle$ is then used for denoising. This technique effectively maintain the features from the first frame in subsequent frames.
@@
FLASH adopts these components in its base video diffusion model, and designs the Motion Alignment Module and the Detail Enhancement Decoder on top of it.
@@
A2.2 Strongly Augmented Videos
To create a strongly augmented version of an original video, we sequentially apply Gaussian blur and random color adjustments to the original video. This process is designed to preserve the original motion while altering the appearance uniformly across all frames.
  • Gaussian blur: A kernel size is randomly selected from a predefined range, as specified in Sec. A3.2. This kernel size is used to apply Gaussian blur to every frame of the original video, ensuring a uniform level of blur throughout.
  • Random color adjustments: After applying Gaussian blur, we randomly adjust the brightness, contrast, saturation, and hue of the video. For each property, an adjustment factor is randomly chosen from its respective predefined range, detailed in Sec. A3.2. The adjustment with the chosen factor is applied uniformly across all frames to maintain consistent color alterations without introducing cross-frame inconsistencies. We implement it with the ColorJitter function in PyTorch.
@@
Multi-scale Detail Propagation. Before feeding the two features $\bm{g}_{l}^{1}$ and $\bm{h}_{l}^{i}$ to the two branches, we first interpolate $\bm{g}_{l}^{1}$ to match the spatial size of $\bm{h}_{l}^{i}$ and use a fully connected layer to adjust $\bm{g}_{l}^{1}$ to the same number of channels as $\bm{h}_{l}^{i}$, resulting $\tilde{\bm{g}}_{k}^{1}$ as the input of the two branches. The network $\mathcal{N}$ in the Warping Branch is a four-layer convolution network that takes the channel-wise concatenation of $\bm{h}_{l}^{i}$ and $\tilde{\bm{g}}_{l}^{1}$ as input and outputs the spatial displacements $(\Delta p,\Delta q)$. In the Patch Attention Branch, before applying the cross-attention layer $\mathcal{A}$, we use a fully connected layer to transform each patch of the two features into a feature vector. The network $\mathcal{M}$, which generates the fusion weights, is a two-layer convolution network, which takes the channel-wise concatenation of $\bm{h}_{l}^{i}$ and $\tilde{\bm{g}}_{l}^{1}$ as input and outputs the fusion weights $\bm{w}^{i}_{l}$.
@@
Distorted Videos. The details of video distortions are as follows: The random Gaussian blur and random color adjustments follow the implementations described in Sec. A2.2. However, the random color adjustments here differ in that they are applied to only 80% of randomly selected regions rather than to all regions. This modification is intentional, as the goal is to create distorted videos with inconsistent color changes that simulate the distortions in latent videos, rather than to maintain consistent color changes as in Sec. A2.2. For random elastic transformations, displacement vectors are generated for all spatial positions based on random offsets sampled from a predefined range (detailed in Sec. A3.2) and are then used to transform each pixel accordingly. We implement it using the ElasticTransform function in PyTorch.
@@
We conduct experiments on 16 actions selected from the HAA500 dataset [9], which contains 500 human-centric atomic actions capturing the precise movements of human, each consisting of 20 short videos. The selected actions include single-person actions (push-up, arm wave, shoot dance, running in place, sprint run, and backflip), human-object interactions (soccer shoot, drinking from a cup, balance beam jump, balance beam spin, canoeing sprint, chopping wood, ice bucket challenge, and basketball hook shot), and human-human interactions (hugging human, face slapping).
@@
Similarity between training videos in the same action class. Videos within the same action class do not share similar visual characteristics, such as scenes, viewing angles, actor positions, or shot types (e.g., close-up or wide shot), as shown in the examples in Figure A3.
Testing images. For each selected action, we use the first frames from the four testing videos as testing images. Additionally, we search online for two human images depicting a person beginning the desired action as additional testing images.
Figure A3: Similarity between training videos in the same action class. The first row presents three videos depicting the action canoeing sprint, and the second row showcases three videos containing the action push-up.
@@
We use AnimteDiff [28] as the base video generative model. We initialize all parameters with pretrained weights of AnimteDiff. The spatial resolution of generated videos is set to $512\times 512$, and the video length is set to 16 frames.
Training of U-Net. We combine features from the first and current frames as keys and values in the spatial cross-frame attention layers. Following [38, 58], we redefine the sampling probability distribution to prioritize earlier denoising stages. In the Motion Alignment Module, we set $\tau$ to 90 and apply motion feature alignment after each temporal attention layer in the U-Net. Inter-frame correspondence alignment is applied to 50% of the cross-frame attention layers, selected randomly. For simplicity, we replace $\bm{Q}$ and $\bm{K}$ of the augmented video with those of the original video when calculating $S$, instead of directly replacing $S$. Gaussian blur is applied with a randomly sampled kernel size between 3 and 10. Random color adjustment modifies brightness, saturation, and contrast by random factors between 0.5 and 1.5, and modifies hue by a random factor between -0.25 and 0.25. Before applying strong augmentations to the original video, we first perform random horizontal flipping and random cropping on the original video. We only train the temporal attention layers, and the key and value matrices of spatial attention layers. The learning rate is set to $5.0\times 10^{-5}$, with training conducted for 20,000 steps.
Training of Detail Enhancement Decoder. The patch size in the Patch Attention Branch is set to 2. For video distortion, Gaussian blur is applied with a random kernel size between 3 and 10. Random color adjustment use random factors for brightness, saturation, and contrast between 0.7 and 1.3, and a random factor for hue between -0.2 and 0.2. For random elastic transformations, displacement strength is randomly sampled from 1 to 20. We only train the newly added layers, with a learning rate of $1.0\times 10^{-4}$ over 10,000 steps.
@@
Figure A4: AMT user study interface.
Figure A5: Qualitative ablation study on different components of FLASH. #1: baseline model, #2: model trained with strongly augmented videos, #3: with motion feature alignment, #4: with inter-frame correspondence alignment, #5: with both alignments, #6: full model with the Motion Alignment Module and the Detail Enhancement Decoder. See Table 2 for the details of each variant.
@@
In line with previous works [91, 92, 29], we use three metrics based on CLIP [67] to assess text alignment, image alignment, and temporal consistency. (1) Text Alignment: We compute the similarity between the visual features of each frame and the textual features of the text prompt, and average the similarities across all frames. (2) Image Alignment: We compute the similarity between the visual features of each frame and the visual features of the provided reference image, and average the similarities across all frames. (3) Temporal Consistency: We calculate the average similarity between the visual features of consecutive frame pairs to obtain the temporal consistency score. We use ViT-L/14 from OpenAI [67] for feature extraction. In these three metrics, higher scores indicate better performance.
Following [94], we utilize Fréchet distance to compare generated and real videos. We use CD-FVD [15] to mitigate content bias in the widely used FVD [83]. We use VideoMAE [82], pretrained on SomethingSomethingV2 [20], for feature extraction and calculate distance between real and generated videos. In this metric, lower distances indicate better performance.
To evaluate the similarity between generated videos and ground-truth videos in the HAA dataset, we calculate the cosine similarity for each pair of the generated and ground-truth videos. (1) Cosine RGB: We extract video features using I3D [3], pretrained on RGB videos, for both the generated and ground truth videos, and calculate cosine similarity for the pair. (2) Cosine Flow: We extract optical flow using RAFT [80] and then use I3D [3], pretrained on optical flow data, to extract video features for cosine similarity calculation. In these two metrics, higher similarities indicate better performance.
@@
We compare FLASH with several baselines: (1) TI2V-Zero [61], a training-free image animation model that injects the appearance of the reference image into a pretrained text-to-video model. We directly test image animation on its checkpoints. (2) SparseCtrl [25], an image animation model that encodes the reference image with a sparse condition encoder and integrates the features into a text-to-video model. It is trained on large-scale video datasets, and we directly test image animation on its checkpoints. (3) PIA [106], an image animation model that incorporates the reference image features into the noised latent video. It is trained on large-scale video datasets, and we directly test image animation on its checkpoints. (4) DynamiCrafter [94], an image animation model that injects the reference image features into generated videos via cross-attention layers and feature concatenation. It is trained on large-scale video datasets, and we directly test image animation on its checkpoints. (5) DreamVideo [90], a customized video generation model which learns target subject and motion using a limited set of samples. We train it to customize motion for each action using the same training videos as FLASH. (6) MotionDirector [109], a customized video generation model which learns target appearance and motion with limited videos. We train its motion adapter with the same training videos as FLASH and its appearance adapter with the testing reference images. Therefore, MotionDirector has access to more data (the testing reference images) than other methods. (7) LAMP [92], a few-shot image animation model which learns motion patterns from a few videos. We train it with the same training videos as FLASH.
@@
Variant # Videos Per Class joint Training   Cosine RGB ($\uparrow$)   Cosine Flow ($\uparrow$)   CD-FVD ($\downarrow$) Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)
#1 16 83.80 68.06 1023.30 22.53 77.10 95.43
#2 16 83.98 70.61 932.92 22.48 76.72 94.91
#5 16 84.46 72.24 906.31 22.52 76.35 95.01
#1 8 82.50 68.13 995.43 22.70 76.05 94.79
#2 8 83.30 70.09 962.82 22.62 74.37 94.40
#5 8 83.40 72.01 943.54 22.66 75.02 94.51
#1 4 81.40 68.02 1050.03 22.22 72.81 94.24
#2 4 81.88 70.15 1045.49 22.60 72.00 93.83
#5 4 82.22 71.83 1031.87 22.46 72.56 94.22
#5 16 84.46 72.24 906.31 22.52 76.35 95.01
#5 16 85.01 72.32 897.05 22.61 77.47 95.39
Table A2: Ablation studies on different values of $\tau$ for motion feature alignment, different values of $p$ for inter-frame correspondence alignment, and the impact of the Warping Branch and Patch Attention Branch in the Detail Enhancement Decoder.
Variant     $\tau$     $p$    Warping Branch    Patch Attention Branch   Cosine RGB ($\uparrow$)   Cosine Flow ($\uparrow$)   CD-FVD ($\downarrow$) Text Alignment ($\uparrow$) Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)
#3  90 - - - 84.44 71.40 920.39 22.64 76.48 95.06
#3 75 - - - 84.38 71.19 904.25 22.58 76.63 95.16
#3 50 - - - 84.30 70.31 934.84 22.57 77.29 95.14
#3 25 - - - 84.71 69.79 930.53 22.33 76.52 94.85
#4 -  1.0 - - 84.22 69.34 914.12 22.50 76.43 94.91
#4 - 0.5 - - 84.32 71.72 938.21 22.70 76.31 94.84
#5 90 0.5 84.46 72.24 906.31 22.52 76.35 95.01
#6 90 0.5 84.63 71.96 918.61 22.54 76.21 95.35
#6 90 0.5 83.32 72.26 888.05 22.71 74.97 95.13
#6 90 0.5 84.51 72.33 908.39 22.77 76.22 95.31
Table A3: Evaluation of the Detail Enhancement Decoder with DINO-V2.
Variant Image Alignment ($\uparrow$) Temporal Consistency ($\uparrow$)
#5 87.02 97.48
#6 87.75 97.85
@@
We conducted a user study on Amazon Mechanical Turk (AMT), where workers were asked to select the best generated video from a set of candidates. For each action, we randomly selected four different reference images and their corresponding generated videos for this user study. The AMT assessment interface, shown in Figure A4, presented workers with the following instructions: “You will see a reference image on the left and eight human action videos on the right, all generated from that reference image and an action description. Please carefully select the one video in each question that: (1) Best matches the action description and displays the action correctly and smoothly. (2) Maintains the overall appearance of the reference image on the left.” The interface also displayed the reference image and action description.
To identify random clicking, each question was paired with a control question that has obvious correct answer. The control question includes a real video of a randomly selected action alongside clearly incorrect ones, such as a static video, a video with shuffled frames, and a video from the same action class that mismatches the reference image. The main question and the control question were randomly shuffled within each question pair, and each pair was evaluated by 10 different workers. Responses from workers who failed the control questions were marked as invalid.
In total, we collected 488 valid responses. The preference rates for different methods are shown in the pie chart in Figure 3 in the main paper. FLASH was preferred in 66% of the valid responses, significantly outperforming the next best choices, DynamiCrafter(13%) and LAMP (11%).
@@
In line with the main paper, Variant #1 serves as the baseline, excluding both the Motion Alignment Module and the Detail Enhancement Decoder. Variant #2 uses strongly augmented videos for training without any alignment technique. Variants #3, #4, and #5 progressively incorporate motion feature alignment, inter-frame correspondence alignment, and both, respectively, on top of Variant #2. Lastly, Variant #6 builds upon Variant #5 by incorporating the Detail Enhancement Decoder.
@@
Applicability with Fewer Training Videos. To assess the few-shot learning capability of the Motion Alignment Module, we conduct experiments using 8 and 4 videos randomly sampled from each action class. The results are shown in Table A1. Across different numbers of training videos per action class, Variant #5 consistently outperforms Variants #1 and #2 on CD-FVD, Cosine-RGB and Cosine-Flow. The results show that the Motion Alignment Module enhances the motion quality of animated videos in different few-shot configurations.